OpenStudy (anonymous):

Simplify 5^8 divided by 5^4 using the laws of exponents. Verify your answer by expanding the powers, then dividing. Show your work.

OpenStudy (anonymous):

2

OpenStudy (anonymous):

you just subtract the exponents example: let n=any number n^6-n^3=n^3 because 6-3=3 Tp verify by expanding and dividing. times the same number over again. example: 10^5=10x10x10x10x10=100000 example: 3^4=3x3x3x3=81.

OpenStudy (mathstudent55):